OpenStudy (anonymous):

Solve the following equations for x. Identify any extraneous solutions. Show all work. -4sqrt6x+5=sqrt53 plz help completely stuck

OpenStudy (anonymous):

\[-4\sqrt{6x+5}\sqrt{53}\]

OpenStudy (anonymous):

sorry wrong problem

OpenStudy (anonymous):

\[\sqrt{6x+5}=\sqrt{53}\]

hartnn (hartnn):

square both sides, what u get ?

OpenStudy (anonymous):

6x+5=2809

hartnn (hartnn):

actually, \(\sqrt{55}^2 = 55\)

OpenStudy (anonymous):

how did you do that?

hartnn (hartnn):

\(\huge \sqrt{a}=a^{1/2}\) \(\huge (\sqrt a)^2= a^{1/2 \times 2}=a^1 = a\)

OpenStudy (anonymous):

ok im a little confused still

hartnn (hartnn):

which step are you confused ? \((\sqrt{anything})^2=anything\)

hartnn (hartnn):

there,anything should be non-negative

OpenStudy (anonymous):

ok i see that

OpenStudy (anonymous):

how did you cime up with 55 though

hartnn (hartnn):

oh, sorry, \(\sqrt{53}^2=53\)

hartnn (hartnn):

so, 6x+5 = 53 is what u get

OpenStudy (anonymous):

oh ok now i understand

hartnn (hartnn):

can u find x from 6x+5 = 53 ?

OpenStudy (anonymous):

x6x=48

OpenStudy (anonymous):

6x=48

OpenStudy (anonymous):

then i divide 6 from both sides and get x=8

hartnn (hartnn):

correct!

OpenStudy (anonymous):

i have one more can you check it for me

hartnn (hartnn):

sure!

OpenStudy (anonymous):

\[\sqrt{x+2}+4=x\]

OpenStudy (anonymous):

\[\sqrt{x+2}=4x\] is that correct

hartnn (hartnn):

not actually, the 4 was added on left so, to bring it on right side, we need to subtract 4 from both sides...

OpenStudy (anonymous):